I have the following problem to solve and cannot figure it out. Please Help!!!
Solve. 3(x-2)-2x = x-6

Answer by checkley71(8403)

3(x-2)-2x=x-6 or 3x-6-2x=x-6 or 3x-2x-x=6-6 or 0=0